# Generative Modelling of Structurally Constrained Graphs

### v1

@@ G.2 Controlled Molecular Generation
In the Table below, we explore the generation of acyclic molecules by picking the absence of cycles as constraining property. Importantly, in contrast to all the experiments in the paper, we do not train with only graphs that verify the property. Instead, we use the two models trained in the unconstrained setting (from previous section): one with edge-absorbing transitions (ConStruct) and another with marginal transitions (Constrained DiGress+). We then sample from these models using the absorbing noise model and the projector for acyclicity. The results are presented in Table 10.
Table 10: Controlled graph diffusion for acyclic molecules. The constrained property used is the absence of cycles. Constrained DiGress+ denotes a model that was trained with a marginal noise model, but where the sampling is performed using the absorbing noise and projector. Both models were trained in the full QM9 dataset. We present the results for the methods we ran over five sampling runs in the format mean ± standard error of the mean. For each run, we generated 10000 molecules, following the protocol from Vignac et al. [77].
@@
We observe that both methods output only acyclic molecules, which is a necessary consequence of the utilization of the projector. As the set of cyclic molecules is a subset of the set of unconstrained molecules, we verify some repetition among the generated samples. This leads to a decrease in the values of uniqueness for both models when compared to the unconstrained setting. Most remarkably, while the validity of the molecules generated by the model trained with the marginal noise model is significantly lower than the observed one for the unconstrained sampling setting, ConStruct preserves its high validity values (even higher than in the unconstrained setting). This result validates the foundation upon which ConStruct is laid: due to the marginal noise model, the forward process distribution does not match the reverse one when using the projector, harming the molecular validity of the generated instances. In contrast, the edge-absorbing noise model of ConStruct allows the forward and reverse processes to stay in distribution.
@@ Appendix H Variants of ConStruct - Performance Analysis and Extensions
In this section, we perform some ablations to ConStruct/DiGress to further analyse its performance and explore methodological extensions to the proposed method.
@@ H.1 Performance Analysis
However, this observation alone does not explain the slight performance gap. To investigate further, we ran DiGress+ with an edge-absorbing noise model but without projector, designating it as DiGress+ [absorbing] in Table 11. We observe a significant decrease in performance with this modification. This suggests that it is the choice of the noise model that is hindering ConStruct’s performance for this dataset. In fact, ConStruct outperforms DiGress+ [absorbing], emphasizing the relevance of the projector step. In any case, we remark that we did not adjust the variance schedule beyond the one proposed by Austin et al. [3], leaving room for potential improvement in this aspect. Nevertheless, we observe performance dataset dependency by applying the same modified model to the planar dataset (again, see table 11), where it exhibits a significantly better average ratio compared to DiGress+. This observation aligns with recent research indicating that there is no clear evidence that an optimal noise model can be deduced a priori from dataset statistics [75].
@@ H.2 A Posteriori Modifications
  • DiGress+ [projection] - we only apply the projector to the final samples generated by the unconstrained method. For example, in the case of the planarity as target property, we could find the maximal planar subgraphs of the generated samples. This method would necessarily provide a more efficient sampling procedure (as we only execute the projector step once).
We provide the results in synthetic graphs for both methods in Table 11. We observe that DiGress+ (rejection) attains great V.U.N. values, as expected. Nevertheless, we analyse its alarming computational inefficiencies in Section D.3. Additionally, we see that DiGress+ (projection) achieves worse performance than ConStruct. We attribute this result to the fact that such a scheme fails to inform the generative model about the constraining condition throughout the reverse process, thus not harnessing the full expressivity of the diffusion model. We also attribute the anomalously good performance of this method for the tree dataset to the optimal properties of the projector in such case (see Theorem 2).
Finally, considering the two extreme cases described above (blocking edges at every forward pass vs a posteriori modifications), we conclude that ConStruct finds itself in a sweet spot in the trade-off between additional computational burden and integration into the diffusion model.
@@ H.3 Likelihood-based constrained generation
We also consider two variants of ConStruct where the projector is no longer independent from the diffusion model. Instead, we use the associated likelihoods to each of the sampled candidate edges at a given time step $t$, $p(e^{t-1}_{ij}|G^{t})$, to define an order by which we add the edges. Therefore, instead of uniformly sampling them at random, we propose two methods of integrating such information:
  • Deterministic: we add the edges with higher $p(e^{t-1}_{ij}|G^{t})$ first.
  • Stochastic: we sample without replacement from the set of candidates edges, where the probability of sampling each of them is proportional to the respective $p(e^{t-1}_{ij}|G^{t})$.
We present the results for these likelihood-based variants of ConStruct in Table 11. As we can observe, for all the analysed datasets, the three ConStruct variants are statistically equivalent in terms of performance. In terms of efficiency, there is no meaningful difference among the three methods: even though the uniformly random sampling does not have to sort the order of the edges nor access the corresponding $p(e^{t-1}_{ij}|G^{t})$, the computational overburden of these operations is negligible. We remark that the theoretical analysis performed in Appendix B also holds for the stochastic likelihood-based variant, but not for the deterministic one, as the latter is not able to select any permutation from the candidate edges with non-zero probability: up to the degenerate cases where different candidate edges have the same $p(e^{t-1}_{ij}\mid G^{t})$, it always selects candidate edges by the same order.
Table 11: Graph generation performance on synthetic graphs. DiGress+, ConStruct are retrieved from Table 1, from which we follow the same experimental protocol. DiGress+ [absorbing] denotes DiGress+ with an edge-absorbing noise model. DiGress+ [rejection] refers to the baseline that rejects the unconstrainedly generated graphs that do not satisfy the constraint and resamples, while DiGress+ [projection] applies the projector on them. ConStruct [model - det] and ConStruct [model - stoch] denote the deterministic and stochastic likelihood-based variants of ConStruct.
@@ I.1.1 Synthetic Datasets
Figure 6: Uncurated set of dataset graphs (top) and generated graphs by ConStruct (bottom) the planar dataset.
@@ I.1.2 Digital Pathology
We provide plots of the sampled graphs from ConStruct for the low TLS content dataset in Figure 9 and for the high TLS content dataset in Figure 10. We also several snapshots throughout the reverse process of ConStruct in Figure 11 to illustrate it as an edge insertion procedure.
@@
Figure 11: Reverse processes for generation of low (top) and high (bottom) TLS content graphs using ConStruct. We start from a graph without any edge on the left ($t=T$) and progressively construct the graph, as a consequence of the absorbing noise model. The node types switch due to the marginal noise model. On the right, we have a fresh new sample ($t=0$). The phenotypes color key is presented in Figure 4.
@@ I.2 Visualizing Intermediate Graphs (Before and After Projector)
Figure 12: Visualizations of intermediate graphs throughout the reverse process. The notation follows the one of the rest of the paper: we obtain $G^{t-1}$ after projecting $\hat{G}^{t-1}$, which in turn is obtained from $G^{t}$ through the diffusion model. From the new edges obtained in $\hat{G}^{t-1}$, we color them in green when they do not break the constraining property and in red otherwise. We can observe that the red edges are rejected. To better emphasize the edge rejection by the projector, we do not use a fully trained model and use a trajectory length, $T$, smaller than usual.

### v2

@@ G.2 Controlled Molecular Generation
We explore the generation of acyclic molecules by picking the absence of cycles as constraining property for ConStruct. Importantly, in contrast to all the experiments in the paper, we do not train with only graphs that verify the property. Instead, we use the two models trained in the unconstrained setting (from previous section): one with edge-absorbing transitions (ConStruct) and another with marginal edge transitions (DiGress+). We then sample from these models using the absorbing noise model and the projector for acyclicity. The results are presented in Table 10.
Table 10: Controlled graph diffusion for acyclic molecules. The constrained property used is acyclicity. Constrained DiGress+ denotes a model that was trained with a marginal noise model, but where the sampling is performed using the edge-absorbing noise model and projector. Both models were trained in the full QM9 dataset. We performed five sampling runs for each method and present their results in the format mean ± standard error of the mean. For each run, we generated 10000 molecules, following the protocol from Vignac et al. [79]. The FCD is computed using the official implementation from Preuer et al. [63].
@@
We observe that both methods output only acyclic molecules, which is a necessary consequence of the utilization of the projector. As the set of acyclic molecules is a subset of the set of unconstrained molecules, we verify some repetition among the generated samples. This leads to a decrease in the values of uniqueness for both models when compared to the unconstrained setting. Most remarkably, while the validity of the molecules generated by the model trained with the marginal noise model is significantly lower than the observed one for the unconstrained sampling setting, ConStruct preserves its high validity values (even higher than in the unconstrained setting). This result validates the foundation upon which ConStruct is laid: with the marginal noise model, the forward process distribution does not match the reverse one when employing the projector, harming the molecular validity of the generated instances. In contrast, the edge-absorbing noise model of ConStruct allows the forward and reverse processes to match, staying in distribution.
@@ Appendix H Variants of ConStruct - Performance Analysis and Extensions
In this section, we perform some ablations to ConStruct/DiGress+ to further analyse its performance and explore methodological extensions to the proposed method.
@@ H.1 Performance Analysis
However, this observation alone does not explain the slight performance gap. To investigate further, we ran DiGress+ with an edge-absorbing noise model but without projector, designating it as DiGress+ [absorbing] in Table 11. We observe a significant decrease in performance with this modification. This suggests that it is the choice of the noise model that is hindering ConStruct’s performance for this dataset. In fact, ConStruct outperforms DiGress+ [absorbing], emphasizing the relevance of the projector step. In any case, we remark that we did not adjust the variance schedule beyond the one proposed by Austin et al. [4], leaving room for potential improvement in this aspect. Importantly, we observe a dataset dependency of performance by applying the same modified model to the planar dataset (again, see Table 11), where it exhibits a significantly better average ratio compared to DiGress+. This observation aligns with recent research indicating that there is no clear evidence that an optimal noise model can be deduced a priori from dataset statistics [77].
@@ H.2 A Posteriori Modifications
  • DiGress+ [projection] - we only apply the projector to the final samples generated by the unconstrained method. For example, in the case of planarity as target property, we could find the maximal planar subgraphs of the generated samples. This method would necessarily provide a more efficient sampling procedure (as we only execute the projector step once).
We provide the results in synthetic graphs for both methods in Table 11. We observe that DiGress+ [rejection] attains great V.U.N. values, as expected. Nevertheless, we analyse its alarming computational inefficiencies in Section D.3. Additionally, we see that DiGress+ (projection) achieves worse performance than ConStruct. We attribute this result to the fact that such a scheme fails to inform the generative model about the constraining condition throughout the reverse process, thus not harnessing the full expressivity of the diffusion model. We also attribute the anomalously good performance of this method for the tree dataset to the optimal properties of the projector in such case (see Theorem 2).
Finally, considering the two extreme cases described above (blocking edges at every forward pass vs a posteriori modifications), we conclude that ConStruct finds itself in a sweet spot in the trade-off between additional computational burden and constraint integration into the generative process.
@@ H.3 Likelihood-based constrained generation
We also consider two variants of ConStruct where the projector is no longer independent from the diffusion model. Instead, we use the associated likelihoods to each of the sampled candidate edges at a given time step $t$, $p_{\theta}(e^{t-1}_{ij}|G^{t})$, to define an order by which we add the edges. Therefore, instead of uniformly sampling them at random, we propose two methods of integrating such information:
  • Deterministic: we add the edges with higher $p_{\theta}(e^{t-1}_{ij}|G^{t})$ first.
  • Stochastic: we sample without replacement from the set of candidates edges, where the probability of sampling each of them is proportional to the respective $p_{\theta}(e^{t-1}_{ij}|G^{t})$.
We present the results for these likelihood-based variants of ConStruct in Table 11. As we can observe, for all the analysed datasets, the three ConStruct variants are statistically equivalent in terms of performance. In terms of efficiency, there is no meaningful difference among the three methods: even though the uniformly random sampling does not have to access $p(e^{t-1}_{ij}|G^{t})$, nor sort the order of the corresponding edges, the computational overburden of these operations is negligible. We remark that the theoretical analysis performed in Appendix B also holds for the stochastic likelihood-based variant, but not for the deterministic one, as the latter is not able to select any permutation from the candidate edges with non-zero probability: up to the degenerate cases where different candidate edges have the same $p(e^{t-1}_{ij}\mid G^{t})$, it always selects candidate edges by the same order.
Table 11: Graph generation performance on synthetic graphs. DiGress+, ConStruct are retrieved from Table 1, from which we follow the same experimental protocol. DiGress+ [absorbing] denotes DiGress+ with an edge-absorbing noise model. DiGress+ [rejection] refers to the baseline that rejects the unconstrainedly generated graphs that do not satisfy the constraint and resamples until the intended number of valid graphs is reached, while DiGress+ [projection] directly applies the projector on unconstrainedly generated graphs. ConStruct [model - det] and ConStruct [model - stoch] denote the deterministic and stochastic likelihood-based variants of ConStruct.
@@ I.1.1 Synthetic Datasets
Figure 6: Uncurated set of dataset graphs (top) and generated graphs by ConStruct (bottom) for the planar dataset.
@@ I.1.2 Digital Pathology
We provide plots of the sampled graphs from ConStruct for the low TLS content dataset in Figure 9 and for the high TLS content dataset in Figure 10. We also provide several snapshots throughout the reverse process of ConStruct in Figure 11 to illustrate it as an edge insertion procedure.
@@
Figure 11: Reverse processes for generation of low (top) and high (bottom) TLS content graphs using ConStruct. We start from a graph without any edge on the left ($t=T$) and progressively build the graph, as a consequence of the absorbing noise model. The node types switch along the trajectory due to the marginal noise model. On the right, we have a fresh new sample ($t=0$). The phenotypes color key is presented in Figure 4.
@@ I.2 Visualizing Intermediate Graphs (Before and After Projector)
Figure 12: Visualizations of intermediate graphs throughout the reverse process. The notation follows the one of the rest of the paper: we obtain $G^{t-1}$ after applying the projector on $\hat{G}^{t-1}$, which in turn is obtained from $G^{t}$ through the diffusion model. From the new edges obtained in $\hat{G}^{t-1}$, we color them in green when they do not break the constraining property and in red otherwise. We can observe that the red edges are rejected. To better emphasize the edge rejection by the projector, we do not use a fully trained model and use a trajectory length, $T$, smaller than usual, resulting in less accurate edge predictions.
@@
For the particular case of the digital pathology setting, while the generated graphs are able to mimic clinically relevant structures, they remain too small to have any direct clinical impact. Pathologists use whole-slide images for informed decisions, whose corresponding cell graphs typically comprise a total number of nodes 3 to 4 orders of magnitude above the graphs generated at this stage.